OpenStudy (anonymous):

The equation of a circle is shown below. (x -5)2 + (y + 2)2 = 64 The radius of the circle is ________ units

OpenStudy (anonymous):

there is no pic :[

OpenStudy (anonymous):

8

OpenStudy (anonymous):

cool i thought it was going to be 6 or 9

OpenStudy (angela210793):

@SNSDYoona explain to Tamatchi u u got 8

OpenStudy (anonymous):

ohh ok. I got 8 because i square root 64

OpenStudy (anonymous):

oh ok then i dont think i did that :]

OpenStudy (anonymous):

the circle equation is (x+h)^2 + (y+d)^2 = r^2 h and d are the x and y coordinates and r is the radius of the circle

OpenStudy (angela210793):

nice :) exaplain how u get the asnwer :)

OpenStudy (anonymous):

i no from there

OpenStudy (anonymous):

i didnt sqr root it

OpenStudy (anonymous):

i just wanted to double check my answer

OpenStudy (phi):

You should correct this: the circle equation is (x+h)^2 + (y+d)^2 = r^2 if (h,k) are the center of the circle, the equation should read (x-h)^2 + (y-k)^2 = r^2

OpenStudy (anonymous):

depends if h and k are negative or positive x n y values

OpenStudy (anonymous):

yea i shud have put (x-h)^2 + (y-k)^2 = r^2

OpenStudy (anonymous):

would it give you the same answer

OpenStudy (anonymous):

umm yea

OpenStudy (anonymous):

im very good in science math is m weakness y

OpenStudy (anonymous):

wait u have to square root it o.O

OpenStudy (anonymous):

my